Consider a village well which is operated by a pump with capacity 200 litres a day. The marginal cost of drawing water from the well is 5x if x litres of water has been drawn and if the pump has not crossed its capacity. The Marginal cost jumps to 20x litre if the pump’s capacity for the day is exhausted. The water can be sold in the open market at 40 rupees per litre. There are 50 households in the village.


1
Expert's answer
2021-04-12T13:09:58-0400
"\frac {200}{50}=4"

liters per day per household.


"5\times4=20"

"20\times (200-x)\times40=800(200-x)"